OpenStudy (anonymous):

3y-2x+z-2x+3(x+y) A. -x+4y+z B. -x+6y+z C. x+4y+z D. 3x+4+z

OpenStudy (matt101):

Answer is B

OpenStudy (anonymous):

3y - 2x + z - 2x + 3(x + y) --- distribute through the parenthesis 3y - 2x + z - 2x + 3x + 3y -- now combine like terms -2x - 2x + 3x = -4x + 3x = -x 3y + 3y = 6y and then there is the z put it all together and you get : -x + 6y + z Just thought you might want to see how the answer was found :)